What is the mole fraction of the solvent that contains 10.0 grams of Phosphorus pentachloride that has been dissolved in 750 mL of water?


1
Expert's answer
2022-02-21T17:19:04-0500

750ml = 750g

Molar mass of water= 18.01528g/mol

Molar mass of phosphorus pentachloride= 208.24g/mol

Mole fraction= mole of target substance divided by total moles involved

Mole of water=(750÷18.01528)

= 41.63

Mole of phosphorus pentachloride=(10÷208.24)= 0.048

Mole total=(41.63+0.048)=41.678

Mole fraction of solvent therefore is;

(41.63÷41.678)=0.9988

= 0.9988